Amir wants to build a box to keep his toys in. The box is 6 m long, 2 m wide, and 3 m high. The cost of materials to build this box is $2.00 for each cubic metre.


1
Expert's answer
2021-05-04T13:13:03-0400

Solution. Find the volume of the box


"V=lwh"

where l=6m is the box length; w=2m is the box width; h=3m is the box height.

As result get


"V=6m\times2m\times3m=36m^3"

According to the condition of the problem the price per cubic meter is two dollars. Hence the cost of materials to build this box is equal to


"C=36m^3\times 2\frac{dollars}{m^3}=72 dollars"

Answer. 72 dollars.